OpenStudy (anonymous):

Solve for x (x-2)^-4=625

OpenStudy (4meisu):

Is this a log question?

OpenStudy (anonymous):

no

OpenStudy (4meisu):

(x-2)^4=625 = (x^4 + 16) = 625 x^4 = 609 x= 5

OpenStudy (anonymous):

I got 7 I am really confused

OpenStudy (anonymous):

that (x-2)^4 is -4 not 4

OpenStudy (4meisu):

How did you get 7? What's your working?

OpenStudy (anonymous):

(x-2)=625^(1/4) x-2=5 x=7

OpenStudy (anonymous):

Can someone help me please

OpenStudy (4meisu):

Oh my bad, I forgot to put the negative before the 4. Then yep, 7 is the answer!

OpenStudy (4meisu):

Actually no

OpenStudy (4meisu):

you're supposed to solve what's in the bracket first.

OpenStudy (4meisu):

you can't bring the ^4 to the other side.

OpenStudy (anonymous):

\[(x-2)^{-4}=625\]

OpenStudy (anonymous):

oh ok but then x=7 is correct

OpenStudy (4meisu):

No it would not be. If x = 7 the answer would be 0.016, not 625.

OpenStudy (anonymous):

I am really confused

OpenStudy (4meisu):

You have to first solve what's in the brackets. So (x-2) ^-4

OpenStudy (anonymous):

ok got that then I have x^-4 then what

OpenStudy (anonymous):

this \((x-2)^{-4}\) means this \(\frac{1}{(x-2)^4}\)

OpenStudy (anonymous):

you have \[\frac{1}{(x-2)^4}=625\] so \[(x-2)^4=\frac{1}{625}\]

OpenStudy (anonymous):

Yes^, but this means there are real and complex answers.?

OpenStudy (anonymous):

that means \[x-2=\pm\frac{1}{5}\]

OpenStudy (anonymous):

if \[x-2=-\frac{1}{5}\] then \[x=-\frac{1}{5}+2=\frac{9}{5}\] and if \[x-2=\frac{1}{5}\] then \[x=\frac{1}{5}+2=\frac{11}{5}\]

OpenStudy (anonymous):

so your two answers are \[x=\frac{9}{5}\] or \[x=\frac{11}{5}\]

OpenStudy (anonymous):

and 2 - i/5 2 + i/5 if you want the complex parts